# A Langevin sampling algorithm inspired by the Adam optimizer

### v1

@@ 5.3.4 MNIST Image Classification on a CNN
Since image classifiers usually adopt convolutional neural networks (CNN) rather than fully connected ones, we repeat the experiment from the previous section on a simple CNN using three convolutional layers. The network architecture is given in Appendix E. Unless explicitly restated, the hyperparameters are the same as in the previous section. Fig. 17 (left) shows the results of a single SamAdams trajectory compared to three different BAOAB trajectories, each one corresponding to a different stepsize (the smallest stepsize adopted by SamAdams, the mean stepsize, and the largest stepsize). SamAdams outperforms all BAOAB runs in terms of loss, train, and test accuracy. In particular, the two BAOAB runs at larger stepsizes become unstable during the early epochs and fail to train completely. The smallest BAOAB stepsize leads to reasonable results, but the convergence speed as measured in number of epochs (i.e., compute time) is substantially smaller than for SamAdams, implying enormous computational speed-ups when using the latter. Fig. 17 (right) shows the evolution of $\zeta$ and $\Delta t$ of the SamAdams run compared to the loss. One observes how the algorithm reacts to the instabilities visible in the loss during the early training phase by rapidly reducing the stepsize $\Delta t$, well below the level of the learning rates used by the two unstable BAOAB runs. Only the BAOAB run using the smallest $\Delta t$ adopted by SamAdams remained stable, implying that small stepsizes are necessary to make it through the early stage of training. However, from epoch 5 onwards SamAdams slowly increases its stepsize again until it stabilizes at $\sim 0.0011$, more than three times the size of the stepsize used by the stable BAOAB run. This explains SamAdams faster convergence in loss and accuracy. From looking at the loss curves and the obtained $\Delta t$ by SamAdams, it seems like the trajectories start on a plateau (allowing for a rapid increase in $\Delta t$ at the very beginning), then descend down through an irregular landscape (leading to rapid damping in $\Delta t$ and breakdown of two of the three BAOAB runs), then reach a widened basin (allowing for moderate increase of $\Delta t$ again). This challenges the conventional wisdom in deep learning to use large learning rates at the start of training and successive decrease of learning rates during later phases [24, 4, 26]. Potentially, one could use adaptive stepsize schemes in the place of conventional learning rate schedulers, an idea the exploration of which we leave for future work.
@@
Figure 17: Training of a CNN on MNIST. Left: Train loss, train and test accuracies. BAOAB was run at three different stepsizes: the smallest, largest, and mean stepsize used by SamAdams. Right: SamAdams results for loss (same as on the left), $\zeta$, and $\Delta t$. The dashed lines correspond to the stepsizes used by BAOAB. Hyperparameters: $\Delta\tau=0.002$, $\alpha=500$.
We now examine accuracies obtained by the two schemes via posterior sampling and averaging. Similar to the previous section, we run 100 independent trajectories, initialized as before. The hyperparameters are the same as in Fig. 17. Each trajectory is run for 60 epochs where the train and test accuracies are computed every epoch after the first 40 epochs. For each trajectory, we time-average the obtained samples and then average the results across trajectories. The SamAdams runs are executed first as the BAOAB stepsizes will be obtained from the adaptive stepsize values $\Delta t$ used by SamAdams (pooled from all trajectories). BAOAB is run with three different learning rates [^7] We use $h$ here to denote BAOAB’s stepsize to avoid confusion. $h$ (i.e., 300 BAOAB runs in total): The mean of the pooled $\Delta t$ (denoted by $h=\langle\Delta t\rangle$), the mean of the smallest 10% of the pooled $\Delta t$ (denoted by $h=\langle\Delta t\rangle_{\text{small}}$), and the mean of the largest 10% ($h=\langle\Delta t\rangle_{\text{large}}$). For reference, these values are $\langle\Delta t\rangle_{\text{small}}=0.00046$, $\langle\Delta t\rangle=0.00093$, and $\langle\Delta t\rangle_{\text{large}}=0.0014$. Note that this way of choosing the stepsizes of BAOAB is different from the last section where we took the point of view that a fair comparison between adaptive-stepsize and constant-stepsize schemes is made by considering the first 10% of obtained $\Delta t$ rather than the largest or smallest 10%. The results together with 95% confidence intervals are given in Table 2.
SamAdams BAOAB $h=\langle\Delta t\rangle_{\text{small}}$ BAOAB $h=\langle\Delta t\rangle$ BAOAB $h=\langle\Delta t\rangle_{\text{large}}$
Train Accuracy (%) 98.48$\pm$0.07 70.07$\pm$7.31 11.65$\pm$1.97 10.26$\pm$0.52
Test Accuracy (%) 97.97$\pm$0.05 70.26$\pm$7.33 11.63$\pm$1.98 10.23$\pm$0.53
Table 2: Mean accuracies and 95%-confidence intervals obtained by sampling the posterior of a CNN on the MNIST dataset. 100 independent trajectories were run per column. The three BAOAB stepsizes were obtained from the adaptive stepsize values used by SamAdams (see main text). The other hyperparameters are as in Fig. 17.
SamAdams reaches high accuracies with small variance, i.e. high reliability. Comparable to the single-trajectory results in Fig. 17, the two larger BAOAB stepsizes fail almost completely (since there are 10 classes, 10% corresponds to random class label assignment). Only when run at the smallest of the three chosen stepsizes does BAOAB train properly, but with far worse results than SamAdams. The fact that this stepsize only leads to ${}70\%$ accuracies when a similar stepsize led to accuracies of more than 95% in Fig. 17 implies that the variance at this stepsize must be high (also evident by the large confidence intervals in the table). In fact, when plotting the histograms of the (time-averaged) train accuracies of SamAdams and the $\langle\Delta t\rangle_{\text{small}}$-BAOAB runs, we see that while the SamAdams accuracies all lie north of 95%, BAOAB yields a significant number of accuracies close to 10%, see Fig. 18.
@@
Figure 18: Histograms of time-averaged train accuracies for 100 independent trajectories of SamAdams (left) and BAOAB (right) corresponding to the results in Table 2. BAOAB was run at $h=\langle\Delta t\rangle_{\text{small}}=0.00046$, see main text for explanation.
@@ 6 Conclusion
The method can be adapted to large-scale Bayesian machine learning, and is likely to show advantages in relation to models that are typically currently trained using Adam and its derivatives in the deep learning context [80], whether for natural language processing [78], diffusion models [51] or some other type of machine learning application. As sampling and BNN frameworks are introduced to address a wider range of AI challenges, we expect algorithms such as the one described here will be in high demand. Manual adjustment of the stepsize (learning rate scheduling) is often used in machine learning applications; the flexible nature of our framework suggests the possibility of an automated approach which can simplify training workflows if the right choice of monitor function and other aspects can be identified (which may not always be as simple as the norm of the gradient). The strong relation between the stepsize used in training, the batch size (amount of gradient noise) and the generalization error makes the new method interesting for active learning settings in which batch sizes vary in time [2]. Although it is not the main target of this work, we believe that SamAdams (or a similar method based on the adaptation framework presented in this article) might be of interest to computational scientists simulating physical models in which forces may increase preciptously during integration, often requiring the use of small stepsizes compared to the long simulation times that have to be realized.
@@ Acknowledgements
This research was supported by the MAC-MIGS Centre for Doctoral Training (grant EPSRC EP/S023291/1). The authors wish to acknowledge Katerina Karoni for helpful discussions at an early stage of the project and Daniel Paulin for advice on the neural network studies. We also thank Gilles Vilmart for pointing out the article [99] which we have used to establish the weak convergence of our method and Michael Tretyakov for a helpful discussion on the numerical stability of SDE discretizations.
@@ Appendix A General Form of the SamAdams Algorithm
${\rm d}x_{\tau}$ $=$ $\psi(\zeta)a(x_{\tau},t(\tau)){\rm d}{\tau}+\sigma(x_{\tau},t(\tau))\sqrt{\psi(\zeta)}{\rm d}W_{\tau},$ (27)
${\rm d}\zeta$ $=$ $f(x_{\tau},\zeta(\tau)){\rm d}\tau,$ (28)
${\rm d}t$ $=$ $\psi(\zeta){\rm d}\tau.$ (29)
@@ Appendix C Further Details of Stepsize Control Mechanism
Appendix D Example with Gradient Noise in Logistic Regression
@@
As the batch size decreases, the accuracy in both methods drops. However, due to the gradient noise, the otherwise simple dynamics (logistic regression posterior is convex) becomes more unstable, leading to automatic stepsize reduction in SamAdams. The sampling bias introduced by gradient noise from data subsampling scales with the stepsize, so a reduction of stepsize for decreasing batch size can restore accuracy. Note that the performance of BAOAB can be restored by picking a smaller stepsize as well. We merely wish to highlight the benefit of SamAdams to automatically react to induced gradient noise. Fig. 23 shows the batch size-dependent $\Delta t$-histograms obtained by SamAdams.
Figure 23: SamAdams (ZBAOABZ) $\Delta t$ histograms against batch size for the experiment of Fig. 22.
@@
The stepsize adaptation is able to change $\Delta t$ dynamically and almost instantaneously in accordance with the batch size. While we did not find significant differences in the resulting accuracies of SamAdams and BAOAB when using dynamically changing batch sizes, we believe that this is a consequence of the simplicity of the problem. Logistic regression on comparably small datasets does not require stochastic gradients, neither for efficiency nor to escape local minima. It would be interesting to examine the combined impact of stepsize adaptation and gradient noise on large-scale deep learning models, where gradient subsampling is quintessential for efficient training, or in active learning settings, in which the batch size of different training iterations is allowed to vary.
Appendix E CNN Architecture

### v2

@@ 5.3.3 MNIST Image Classification on a CNN
As image classifiers usually adopt convolutional neural networks (CNN) rather than fully-connected ones, we repeat the experiment from the previous section on a simple CNN using three convolutional layers. The network architecture is given in Appendix G. Unless explicitly restated, the hyperparameters are the same as in the previous section.
Fig. 16 (left) motivates the use of variable stepsizes. It shows the results of a single SamAdams trajectory compared to three different BAOAB trajectories, each one corresponding to a different stepsize: (i) the smallest stepsize adopted by SamAdams, (ii) the mean stepsize, and (iii) the largest stepsize. This way of choosing the stepsizes of BAOAB to compare it to SamAdams is different from that used in the last section, where we took the point of view that a fair comparison between adaptive-stepsize and constant-stepsize schemes is made by considering the first 10% of the obtained $\Delta t$ as well as the mean of $\Delta t$.
@@
Figure 16: Training of a CNN on MNIST. Left: Train loss, train- and test accuracies. BAOAB was run at three different stepsizes: the smallest, largest, and mean stepsize used by SamAdams. Right: SamAdams results for loss (same as on the left), $\zeta$, and $\Delta t$. The dashed lines correspond to the stepsizes used by BAOAB. Hyperparameters: $\Delta\tau=0.002$, $\alpha=500$.
The two larger BAOAB stepsizes fail completely, while the smallest stepsize leads to reasonable results, but with a convergence speed (as measured in number of epochs, i.e., compute time) substantially reduced in comparison to SamAdams, implying enormous computational speed-ups when using the latter. Fig. 16 (right) shows the evolution of $\zeta$ and $\Delta t$ of the SamAdams run compared to the loss. One observes that the algorithm reacts to the instabilities visible in the loss during the early training phase by rapidly reducing the stepsize $\Delta t$, well below the level of the learning rates used by the two unstable BAOAB runs. Only the BAOAB run using the smallest $\Delta t$ adopted by SamAdams remained stable, implying that small stepsizes are necessary to make it through the early stage of training. However, from epoch 5 onward SamAdams slowly increases its stepsize again, until it stabilizes at $\sim 0.0011$, more than three times the size of the stepsize used by the stable BAOAB run. This explains SamAdams faster convergence in loss and accuracy. Looking at the loss curves and the obtained $\Delta t$ by SamAdams, it appears that the trajectories start on a plateau (allowing for a rapid increase in $\Delta t$ at the very beginning), then descend through an irregular landscape (leading to rapid damping in $\Delta t$ and breakdown of two of the three BAOAB runs), before reaching a widened basin (allowing for moderate increase of $\Delta t$ again). This challenges the conventional wisdom in deep learning to use large learning rates at the start of training and to decrease of learning rates during later phases [25, 4, 27]. Potentially, one could use adaptive stepsize schemes in place of conventional learning rate schedulers; we leave the detailed study of this intriguing possibility for future work.
[…]
Figure 17: Training of a CNN on MNIST. The top row shows SamAdams results. The bottom three rows show BAOAB results for different stepsizes. Each figure shows 100 trajectories. Left: Loss vs. dynamics time $t$. Middle: (Test) accuracy vs. dynamics time $t$. Right: (Test) accuracy vs. epochs.
The first two columns in Fig. 17 show that all trajectories of both samplers and all examined stepsizes experience instabilities during the early stages of training, at $t\approx 0.03$. This shows that the trajectories indeed follow similar paths [^8] To be more precise, they encounter similar landscape topologies at similar times. They don’t actually follow similar paths as they are initialized randomly and independently., roughly independently of the stepsize. Naturally, this independency will only hold for a certain stepsize range. The BAOAB runs in particular become unstable; while many of them are able to recover from this and begin to converge to high accuracies, some trajectories fail completely. The fraction of trajectories that are able to recover and their convergence speed both decrease with increasing stepsize. In contrast, SamAdams is able to mitigate these issues as evident by the much smaller spikes in the curves, and all trajectories converge to high accuracies, barely being affected by the instabilities. This is particularly impressive given its mean stepsize being twice as large as the largest tested BAOAB stepsize. In order for BAOAB to traverse the unstable early-training region as unobstructed as SamAdams, it would require a smaller stepsize than the ones used in the figure, which would come at the cost of needing to execute many more iterations/epochs. The plots against epoch demonstrate the efficiency gains from using SamAdams, which provides high accuracy as early as epoch 20.
We now compute posterior averages using the samples obtained from the trajectories. Fig. 18 shows the accuracies of the different samplers and stepsizes at certain training epochs. At each such epoch, a time average is performed over the previous 10 epochs (where each epoch yields one sample), followed by the average over the 100 trajectories. The shaded areas denote 95%-confidence intervals (based on t-statistics).
Figure 18: Test accuracies of a CNN on MNIST obtained through Bayesian sampling. Every 10 epochs, the samples of the last 10 epochs were averaged (time average) and the results were averaged across 100 independent trajectories for each of the curves. The shaded areas denote 95%-confidence intervals.
We see that SamAdams converges to high accuracies much faster than all BAOAB runs, none of which reach comparable accuracies even at the end of training. Furthermore, due to several BAOAB trajectories failing (as shown in Fig. 17) the variance of the BAOAB results are much larger than those of SamAdams; the extremely narrow confidence interval, the shaded area behind the blue curve, is almost too narrow to see. The final accuracies and confidences are shown by Table 3.
Sampler, Stepsize Test Accuracy (%)
SamAdams, $\langle\Delta t\rangle=0.00092$ $\textbf{97.90}\pm\textbf{0.06}$
BAOAB, $h=0.00046$ $73.95\pm 6.96$
BAOAB, $h=0.0003$ $92.24\pm 2.71$
BAOAB, $h=0.0002$ $82.11\pm 2.87$
Table 3: Mean accuracies and 95%-confidence intervals obtained by sampling the posterior of a CNN on the MNIST dataset. Values correspond to the epoch 60 results in Fig. 18.
The particularly large variance for BAOAB at $h=0.00046$ comes from the large number of trajectories that fail as a consequence of the early-stage instabilities, see again Fig. 17 and also Fig. 19 which shows the histograms of obtained accuracies after 60 epochs for that BAOAB setting compared to SamAdams.
These experiments demonstrate that the enhanced stability and computational efficiency of SamAdams observed on the artificial planar problems of the previous sections extend to complex models on real-world data, and even establish that the mechanism is similar.
We note that the benefits of the adaptive stepsizes when sampling the posterior of the CNN of this section mainly stem from the early stages of training, during which the sampler equilibrates from the random initial conditions to lower regions of the loss landscape. There are alternative approaches to BNN sampling. For example, one can perform BNN experiments by first running an optimizer to find a local mode, then running an MCMC scheme starting from within that mode to sample the mode vicinity. This can either be done by adding a confining potential as in [76] or by using low temperatures or stepsizes [112, 111]. Our adaptive stepsize framework can easily be used in conjunction with the ideas in those works. In our experiments, from looking at the comparably small fluctuation of $\zeta$ and $\Delta t$ in the later stages of training (see Fig. 16), the necessity for adaptive stepsizes in the mode vicinity is much smaller than during the early stages of equilibration. It is easily conceivable, however, that more complex models and datasets lead to basin topologies which would benefit from adaptive stepsizes as well. We leave the exploration of larger models and different approaches to BNN sampling with adaptive stepsizes for future work.
@@
Figure 19: Histograms of time-averaged train accuracies for 100 independent trajectories of SamAdams (red) and BAOAB (blue) corresponding to the results in Table 3. BAOAB was run at $h=0.00046$.
@@ 6 Conclusion
The method can be adapted to large-scale Bayesian machine learning, and is likely to show advantages in relation to models that are typically trained using Adam and its derivatives in the deep learning context [82], whether for natural language processing [80], diffusion models [52] or some other type of machine learning application. As sampling and BNN frameworks are introduced to address a wider range of AI challenges, we expect algorithms such as the one described here will be in high demand. Manual adjustment of the stepsize (learning rate scheduling) is often used in machine learning applications [25, 4, 27]; the flexible nature of our framework suggests the possibility of an automated approach which can simplify training workflows if the right choice of monitor function and other aspects can be identified (which may not always be as simple as the norm of the gradient). The strong relation between the stepsize used in training, the batch size (amount of gradient noise), and the generalization error makes the new method interesting for settings with batch sizes that vary in time, such as active learning [2]. Although it is not the main target of this work, we believe that SamAdams (or a similar method based on the adaptation framework presented in this article) might be of interest to computational scientists simulating physical models in which forces may increase preciptously during integration, often requiring the use of small stepsizes compared to the long simulation times that have to be realized.
@@ Acknowledgements
This research was supported by the MAC-MIGS Centre for Doctoral Training (grant EPSRC EP/S023291/1). The authors wish to acknowledge Katerina Karoni for helpful discussions at an early stage of the project and Daniel Paulin for advice on the neural network studies. We thank Gilles Vilmart for generous assistance in establishing the weak convergence of our method and Michael Tretyakov for helpful input on numerical stability of SDE discretizations.
@@ Appendix A General Form of the SamAdams Algorithm
${\rm d}x_{\tau}$ $=$ $\psi(\zeta_{\tau})a(x_{\tau},t(\tau)){\rm d}{\tau}+\sigma(x_{\tau},t(\tau))\sqrt{\psi(\zeta_{\tau})}{\rm d}W_{\tau},$ (31)
${\rm d}\zeta_{\tau}$ $=$ $f(x_{\tau},\zeta(\tau)){\rm d}\tau,$ (32)
${\rm d}t$ $=$ $\psi(\zeta_{\tau}){\rm d}\tau.$ (33)
@@ Appendix C Further Details of Stepsize Control Mechanism
Appendix D Symmetric Z-Steps
Algorithm 3.1 in the main text consists of a symmetric integrator of the time-rescaled Langevin dynamics (10)-(13), named ZBAOABZ, that adds two Z-steps as defined in Sec. 3 before and after a constant-stepsize LD integrator, here BAOAB. Right after each of the Z-steps, the stepsize $\Delta t$ is updated. The BAOAB iteration is then executed with a stepsize given by the one obtained from the first Z-step. While the symmetry of the LD integrator used, BAOAB, has beneficial impact on the integration error, the symmetric placing of two Z-steps in combination with two evaluations of $\psi(\zeta)$ is less obvious. Since $\hat{\Phi}^{Z}_{\frac{a}{2}}\circ\hat{\Phi}^{Z}_{\frac{a}{2}}=\hat{\Phi}^{Z}_{a}$, the dynamics would not change if the two Z-steps were merged and the Sundman transform $\psi$ was only evaluated once (namely right before the BAOAB iteration to update $\Delta t$). However, as described in Sec. 2.1, the values $\psi(\zeta)$ are needed to reweight the collected samples of an observable $\phi(x_{n},p_{n})$ in order to obtain averages with respect to the correct measure. It is not obvious which value of $\psi(\zeta)$ should be used to reweight a sample generated by a BAOAB iteration: the value of $\psi(\zeta)$ before the BAOAB iteration, i.e., the value that governs the stepsize used by BAOAB, or the value of $\psi(\zeta)$ after the BAOAB iteration, namely the one that is obtained by updating $\zeta$ through the freshly obtained $(x,p)$-samples. Using two half-steps in $\zeta$ is a compromise between the two approaches which we found to outperform either of the two, see Fig. 23.
Figure 23: Comparison of three different Z-step placings concerning the computation of canonical averages of temperature observables. The symmetric placing, ZBAOABZ, corresponds to Alg. 3.1 in the main text. The results were obtained on the star potential (see main text), and averaged over time and 200 trajectories. Simulation settings: $\beta^{-1}=\gamma=\alpha=1,\ \Delta\tau=0.01,\ g(x,p)=\|\nabla U(x)\|^{2},\ \psi=\psi^{(1)}\text{ with }m=0.1,\ M=10,\ r=0.25$.
Appendix E ZBAOABZ Observed Integration Order
While the integrator ZBAOABZ (see Sec. 3) can be interpreted as solving Langevin Dynamics with variable stepsizes in time $t$, it is a conventional constant-stepsize scheme in time $\tau$ (the process given by (10)-(13)), for which conventional error analysis may be used. As mentioned in the main text and shown in [56], the BAOAB integrator is of weak order 2 in the ergodic limit  [^10] The weak order even becomes 4 when considering configurational observables and working in the high-friction limit, $\gamma\to\infty$.. This means that for a given observable $\phi(q)$, $q:=(x,p)$, and stepsize $\Delta t$, we have
$\lim_{n\to\infty}\Big{|}\mathbb{E}\big{[}\phi\big{(}q_{n}\big{)}\big{]}-\mathbb{E}\big{[}\phi\big{(}q(n\Delta t)\big{)}\big{]}\Big{|}<C\Delta t^{2},$ (69)
[…]
Figure 24: Weak error of ZBAOABZ against stepsize $\Delta\tau$, obtained by averaging over time and independent trajectories. The ground truth of the potential energy was obtained via numerical quadrature. Simulation settings: $\gamma=0.1,\ \beta^{-1}=\alpha=1,\ g(x,p)=\|\nabla U(x)\|^{2},\ \psi=\psi^{(1)}$ with $m=0.1,\ M=10,\ r=0.25$.
Appendix F Example with Gradient Noise in Logistic Regression
@@
As the batch size decreases, the accuracy in both methods drops. However, due to the gradient noise, the otherwise simple dynamics [^11] The logistic regression posterior is convex. becomes more unstable, leading to automatic stepsize reduction in SamAdams. The sampling bias introduced by gradient noise from data subsampling scales with the stepsize, so a reduction of stepsize for decreasing batch size can restore accuracy. Note that the performance of BAOAB can be restored by picking a smaller stepsize as well. We merely wish to highlight the benefit of SamAdams to automatically react to induced gradient noise. Fig. 26 shows the batch size-dependent $\Delta t$-histograms obtained by SamAdams.
Figure 26: SamAdams $\Delta t$ histograms against batch size for the experiment of Fig. 25.
@@
The stepsize adaptation is able to change $\Delta t$ dynamically and almost instantaneously in accordance with the batch size. While we did not find significant differences in the resulting accuracies of SamAdams and BAOAB when using dynamically changing batch sizes, we believe that this is a consequence of the simplicity of the problem. Logistic regression on comparably small datasets does not require stochastic gradients, neither for efficiency nor to escape local minima. It would be interesting to examine the combined impact of stepsize adaptation and gradient noise on large-scale deep learning models, where gradient subsampling is quintessential for efficient training. For example, there have been several recent works in different contexts about the relationship between stepsize and batch size and about dynamically changing batch sizes (see, e.g., [95, 65, 107, 2]). A potential direction of future research could then aim to answer the question whether the ideas in those works would benefit from a batch size sensitive stepsize adaptation.
Appendix G CNN Architecture